# Polar Mounds on Strangeon Stars: the Neutrino Emission from Ultraluminous X-ray Pulsars

### v1

@@ Abstract
Ultraluminous X-ray pulsars (ULXPs) serve as unique astrophysical laboratories, offering critical insights into accretion physics under extreme conditions, such as strong magnetic fields and super-Eddington accretion rates. Additionally, the nature of pulsars, i.e., the equation of state of supranuclear matter, is still a matter of intense debate, basing on either conventional neutron stars or strange stars in the litterateurs. In this work, we investigate accretion columns of ULXPs based on the strangeon-star model, focusing on the thermal mound at the column base. Accounting for Coulomb and strangeness barriers of the strangeon stars, we find that the mound can reach $0.7-0.95\,\rm km$ in height with temperatures above $10^{9}\,\rm K$, enabling substantial neutrino emission via electron-positron annihilation. Heat transport along the strangeon star surface contributes a luminosity of $10^{36}\,\rm erg\,s^{-1}$, independent of the accretion rate. At low accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons dominate the luminosity, while at higher rates ($10^{21}\,\rm g\,s^{-1}$), photon trapping makes neutrino emission the main cooling channel, with total luminosity exceeding photon emission, which saturates near $10^{41}\,\rm erg\,s^{-1}$. Estimating neutrino fluxes at Earth, we find that only the nearest ULXP, Swift J0243.6++6124, could produce a marginally detectable signal, while most extragalactic sources remain well below background levels. These results emphasize the key role of the thermal mound and strangeon star properties in determining accretion luminosities and neutrino emission, offering insights for future modeling and observations of ULXPs.
[…]
@@ 1 Introduction
Ultraluminous X-ray sources (ULXs) are extragalactic, point-like X-ray emitters not associated with active galactic nuclei, exhibiting X-ray luminosities significantly exceeding $L_{\rm X}>10^{39}\,\rm erg\,s^{-1}$ (P. Kaaret et al., 2017; S. N. Fabrika et al., 2021; A. King et al., 2023). Traditionally, ULXs were interpreted as systems powered either by super-Eddington accretion onto stellar-mass black holes (M. C. Begelman et al., 2006; J. Poutanen et al., 2007) or by accretion onto intermediate-mass black holes (E. J. M. Colbert & R. F. Mushotzky, 1999; J. M. Miller, 2006; T. J. Maccarone et al., 2007; F. Koliopanos et al., 2017). However, the discovery of coherent X-ray pulsations in M82 X-2 (M. Bachetti et al., 2014), which exhibits an extraordinary luminosity of $L_{\mathrm{X}}\simeq 1.8\times 10^{40}\,\rm erg\,s^{-1}$} and a pulse period of $1.37\,\rm s$, provided compelling evidence that some ULXs are powered by accreting pulsars rather than black holes. In these systems, matter channeled by the strong magnetic field onto the stellar surface forms hot spots that produce pulsed X-ray emission as the star rotates.
@@
Since the discovery of M82 X-2, numerous other ULX pulsars (ULXPs) have been identified, and their observed properties are summarized in Table 1. Of particular interest is the possible detection of a narrow cyclotron scattering feature at $\sim 4.5\,\rm keV$ in M51 ULX-8 (M. Brightman et al., 2018). Although pulsations have not been detected from this source, the feature has been interpreted as a proton cyclotron line, implying a surface multipole magnetic field as strong as $B\sim 10^{15}\,\rm G$. The X-ray luminosities of ULXPs are typically in the range of $10^{39}-10^{41}\,\rm erg\,s^{-1}$, which is above the Eddington luminosity for accreting pulsars, $L_{\rm Edd}\approx 1.8\times 10^{38}\,(M/M_{\odot})\,\rm erg\,s^{-1}$, where $M$ is the mass of the compact object. This challenges our understanding of accretion physics under extreme magnetic fields, high temperatures, and super-Eddington regimes, making ULXPs unique laboratories for exploring high-energy astrophysical processes.
@@
The origin of the extremely high luminosity observed in ULXPs remains a subject of ongoing debate. Proposed models for ULXPs span a broad range of surface magnetic field strengths. At relatively low fields ($\sim 10^{11}\,\rm G$), the observed luminosities may be explained by strong radiatively driven outflows and geometrical beaming (M. Middleton & A. King, 2016; M. J. Middleton & A. King, 2017; A. King & J.-P. Lasota, 2019; J.-P. Lasota & A. King, 2023). At the other extreme, magnetar-like fields ($\sim 10^{14}\,\rm G$) can suppress radiation pressure (M. M. Basko & R. A. Sunyaev, 1976; H. Tong, 2015; S. Dall’Osso et al., 2016; K. Y. Ekşi et al., 2015) and confine accretion columns near the magnetic poles (A. A. Mushtukov et al., 2015; H. Inoue, 2020). Beyond the field strength itself, several additional factors may substantially influence the accretion dynamics in ULXPs, including complex, non-dipolar magnetic field configurations (G. L. Israel et al., 2017a; S. S. Tsygankov et al., 2018), enhanced neutrino cooling (A. A. Mushtukov et al., 2018; A. Asthana et al., 2023), and photon-bubble instabilities (J. Arons, 1992; M. C. Begelman, 2006; L. Zhang et al., 2021).
@@
On the other hand, three primary mechanisms can contribute to neutrino emission in X-ray binaries: (i) high-energy protons accelerated within the system, such as in jets, relativistic winds, or pulsar magnetospheres. These protons undergo collisions that predominantly produce pions, which then decay into gamma rays and neutrinos (A. Neronov & M. Ribordy, 2009; N. Sahakyan et al., 2014; O. Kosmas et al., 2023; H. R. Christiansen et al., 2006; L. Ducci et al., 2025); (ii) proton-photon interactions in the magnetosphere of the accreting pulsar (W. Bednarek, 2009); (iii) electron-positron pairs of the accretion column of the bright X-ray pulsars, where pair annihilation leads to neutrino emission (A. A. Mushtukov et al., 2018, 2019; A. Asthana et al., 2023).
@@
In previous studies of electron-positron pair annihilation in the accretion columns of bright X-ray pulsars and ULXPs, however, the properties of the thermal mound, as well as the structure and equation of state (EOS) of the dense matter inside pulsars, were often neglected (A. A. Mushtukov et al., 2018; A. Asthana et al., 2023). In fact, the EOS of supranuclear matter inside compact objects remains one of the most challenging problems in both microphysics and astrophysics, hopefully to be solved with the multi-messenger astronomy (J. M. Lattimer & M. Prakash, 2004; L. Baiotti, 2019). Due to the non-perturbative nature of the fundamental strong interaction at low energies, the EOS of dense matter at nuclear densities is still poorly understood, nevertheless, stangeness should play an important role in this energy scale (X. Lai et al., 2023). E. Witten (1984) proposed that the true ground state of dense matter might be quark matter, composed of nearly free $u$, $d$, and $s$ quarks. This concept suggests that compact objects resembling pulsars could be quark stars (QSs) rather than traditional neutron stars (NSs), as outlined in the MIT bag model in the asymptotic freedom regime (C. Alcock et al., 1986). R.-X. Xu (2003) proposed, however, that building units of supranuclear matter could instead be strange quark clusters (renamed “strangeons”), rather than itinerating quarks, due to the fact that the coupling between them is still very strong (X. Y. Lai & R. X. Xu, 2017). These strangeons, formed by bound $u$, $d$, and $s$ quarks, represent a unique state of matter in which quarks condense in position space rather than momentum space. The term strangeon stars (SSs) was coined to describe such compact objects (R. X. Xu & Y. J. Guo, 2017; X. Y. Lai & R. X. Xu, 2017; H.-B. Li et al., 2022; W.-L. Yuan et al., 2025).
@@
Strangeon matter, like strange quark matter, consists of nearly equal numbers of $u$, $d$, and $s$ quarks. However, in contrast to strange quark matter, quarks in strangeon matter are localized inside strangeons due to the strong quark-quark coupling. NSs, QSs, and SSs share similarities but also differ significantly. In SSs, quarks are localized in strangeons, much like neutrons in NSs. However, unlike neutrons, a strangeon can contain more than three valence quarks, restoring light-flavour symmetry. Additionally, SSs are self-bound by the strong force, with their surface matter also composed of strangeons, similar to QSs (R.-X. Xu, 2003).
@@
SSs can explain many key observational phenomena in astrophysics. The EOS of SSs is sufficiently stiff to account for the observed masses of pulsars (P. Demorest et al., 2010; J. Antoniadis et al., 2013), while pulsar glitches may be attributed to starquakes (E. P. Zhou et al., 2014). Furthermore, SSs are capable of explaining X-ray flares and bursts in magnetar candidates (R.-X. Xu et al., 2006), the plateau phase in gamma-ray bursts (S. Dai et al., 2011), the quasi-periodic oscillations in SGR 1806-20 (H.-B. Li et al., 2023), and shaping the energy budget of various power sources in the context of magnetar starquake triggering mechanisms (W.-Y. Wang et al., 2024). Additionally, the strangeness barrier plays a crucial role in understanding Type I X-ray bursters (Z. Li et al., 2015) and X-ray-dim isolated neutron stars (W. Wang et al., 2017; W.-Y. Wang et al., 2018).
@@
In this work, we investigate the thermal mound properties within the SS framework (R.-X. Xu, 2003) and estimate the neutrino luminosity relevant to ULXPs. A notable feature of the polar mound of SSs is the sharp discontinuity in mass density, resulting from both the Coulomb (R. X. Xu & G. J. Qiao, 1999; J. Hu & R. X. Xu, 2002; R.-X. Xu, 2014) and the strangeness (R.-X. Xu, 2014; Z. Li et al., 2015; W. Wang et al., 2017; W.-Y. Wang et al., 2018) barriers. Most of the total gravitational energy of accretion is stored in the matter above the surface of density discontinuity on the polar cap of an SS, whereas an NS does not exhibit such a discontinuity. This therefore results in a lower effective heat capacity for SS and a higher one for NS, and consequently a high temperature and large neutrino emissivity of SS. The detailed quantitative calculations in the following sections demonstrate consistency with these qualitative results.
@@
  • [1][{1}] M. Bachetti et al. (2014);
  • [2][{2}] G. L. Israel et al. (2017a);
  • [3][{3}] S. Carpano et al. (2018);
  • [4][{4}] R. Sathyaprakash et al. (2019);
  • [5][{5}] S. Trudolyubov et al. (2007);
  • [6][{6}] C. Motch et al. (2013); F. Fürst et al. (2016); G. L. Israel et al. (2017b); F. Fuerst et al. (2018);
  • [7][{7}] E. Quintin et al. (2021);
  • [8][{8}] F. Pintore et al. (2021, 2025);
  • [9][{9}] C. A. Wilson-Hodge et al. (2018);
  • [10][{10}] S.-S. Weng et al. (2017);
  • [11][{11}] A. D. Chandra et al. (2020);
  • [12][{12}] G. A. Rodr´ıguez Castillo et al. (2020);
  • [13][{13}] M. Brightman et al. (2018).
@@ 2.1 Dynamical structure of the accretion column
Figure 1: Schematic depiction of the magnetic polar regions with radius $R_{\rm s}\simeq 10^{5}\,\rm cm$ and height $H_{\rm mound}$ of polar mound in an X-ray pulsar. $\Theta$ is the opening angle of the magnetic funnel. This diagram is adapted from H. Inoue (2020).
@@
Radiation pressure plays a crucial role in determining the dynamical structure of accretion flows in high-luminosity X-ray pulsars. At sufficiently high accretion rates, a shock forms above the stellar surface (H. Inoue, 1975; M. M. Basko & R. A. Sunyaev, 1976; H. Inoue, 2020). This shock must be radiative in nature to effectively dissipate the kinetic energy of the infalling matter (P. A. Becker, 1998). Models of such accretion columns are essential for interpreting observed radiation luminosities (A. A. Mushtukov et al., 2015), pulse profiles (D. Klochkov et al., 2008; P. A. Becker et al., 2012), continuum spectra (P. A. Becker & M. T. Wolff, 2006), cyclotron line spectra (R. Staubert et al., 2019), and polarization observations of high-luminosity X-ray pulsars in high-mass X-ray binaries. Moreover, the accretion column exhibit quasi-periodic oscillations using the magnetohydrodynamics simulations (L. Zhang et al., 2022, 2023, 2025)
@@
In highly magnetized accreting pulsars, the plasma is funneled by the strong dipolar magnetic field toward the magnetic polar caps, forming narrow accretion columns above the stellar surface. In these regions, the inflowing matter is nearly constrained to move along magnetic field lines, and the lateral motion can be neglected compared to the radial inflow. Assuming one-dimensional, steady-state accretion, the structure of the accretion column is determined by the conservation of mass, momentum, and energy (H. Inoue, 1975; M. M. Basko & R. A. Sunyaev, 1976; P. A. Becker, 1998).
Following the variational derivation of H. Inoue (1975), the dynamical equations describing the vertical structure of the magnetic polar regions can be written as (see also H. Inoue, 2020),
@@
$r_{\rm D}=1.4\times 10^{5}\left(\frac{\dot{M}}{10^{17}\rm{g\;s}^{-1}}\right)\;\rm{cm}\,$, (3)
where $\dot{M}$ is the mass accretion rate onto the stellar surface. Eq. (1) describes the evolution of the energy density, balancing gravitational acceleration $(GM/r^{2})$ and radiative dissipation. Eq. (2) governs the density stratification, where the compressive effects of gravity compete with radiative pressure support. Together, these equations determine the vertical structure of the accretion column and provide the foundation for understanding the formation of the thermal mound and the associated energy transport in the polar regions.
We adopt the same boundary conditions as in H. Inoue (2020). At the top boundary of the polar cone, we set the mass density to
@@
In Fig. 2, we show the radiation energy, mass density, pressure, velocity, optical depth, and temperature as functions of the column height for different mass accretion rates. It can be observed that an extremely high mass accretion rate results in a taller accretion column. Moreover, the pressure and temperature at the base of the column remain independent of the accretion rate because the magnetic pressure counterbalances the radiation pressure.
In particular, the temperature in the polar mound exceeds $10^{9}\,\rm K$, leading to energy losses through neutrino emission (G. Beaudet et al., 1967; A. D. Kaminker et al., 1994). According to the results of G. Beaudet et al. (1967), the primary mechanism for neutrino production is the annihilation of electron-positron pairs, resulting in the creation of neutrino–anti-neutrino pairs.
Figure 2: Profiles of physical quantities within the magnetic polar regions as functions of height, with different colors indicating different mass accretion rates. We adopt a stellar model with $M=1.4\,M_{\odot}$ and $R=10^{6}\,\rm cm$. The first column presents the radiation energy density, mass density, and pressure, while the second column shows the corresponding velocity, optical depth, and temperature profiles. Normalization parameters are defined as $\varepsilon_{0}=GM/R$, $\rho_{0}=3B^{2}R/(8\pi GM)\sim 10^{2}\,\rm g\,cm^{-3}$, and $P_{0}=B^{2}/(8\pi)$, assuming a surface magnetic field strength of $B=10^{12}\,\rm G$. The characteristic temperature, $T_{\rm C}$, is calculated via $T_{\rm C}=(\rho\varepsilon/a)^{1/4}$, where $a$ is the radiation constant.

### v2

@@ Abstract
Ultraluminous X-ray pulsars (ULXPs) serve as unique astrophysical laboratories, offering critical insights into accretion physics under extreme conditions, such as strong magnetic fields and super-Eddington accretion rates. Additionally, the nature of pulsars, i.e., the equation of state of supranuclear matter, is still a matter of intense debate, basing on either conventional neutron stars or strange stars. In this work, in order to differentiate the conjectured states of matter, we investigate accretion columns of ULXPs based on the strangeon-star (SS) model, focusing on the thermal mound at the column base. Accounting for Coulomb and strangeness barriers of SSs, we find that the mound can reach $0.7-0.95\,\rm km$ in height with temperatures above $10^{9}\,\rm K$, enabling substantial neutrino emission via electron-positron annihilation. At low accretion rates ($10^{20}\,\rm g\,s^{-1}$), photons dominate the luminosity, while at higher rates ($10^{21}\,\rm g\,s^{-1}$), photon trapping makes neutrino emission the main cooling channel, with total luminosity exceeding photon emission, which saturates near $10^{41}\,\rm erg\,s^{-1}$. Even though the predicted neutrino flux from the nearest system, Swift J0243.6++6124, lies well below the diffuse MeV background–implying that detectable emission would require substantially closer or more luminous sources–these results demonstrate the key role of the thermal mound and SS properties in accretion, providing a foundation for future ULXP studies and suggesting that neutrino observations could, in principle, offer a novel probe of SSs and extreme supranuclear matter.
keywords:
stars: neutron – pulsars: general – X-rays: binaries – neutrinos – dense matter.
pubyear: 2026pagerange: Polar Mounds on Strangeon Stars: the Neutrino Emission from Ultraluminous X-ray PulsarsPolar Mounds on Strangeon Stars: the Neutrino Emission from Ultraluminous X-ray Pulsars
@@ 1 Introduction
Ultraluminous X-ray sources (ULXs) are extragalactic, point-like X-ray emitters not associated with active galactic nuclei, exhibiting X-ray luminosities significantly exceeding $L_{\rm X}>10^{39}\,\rm erg\,s^{-1}$ (Kaaret et al., 2017; Fabrika et al., 2021; King et al., 2023). Traditionally, ULXs were interpreted as systems powered either by super-Eddington accretion onto stellar-mass black holes (Begelman et al., 2006; Poutanen et al., 2007) or by accretion onto intermediate-mass black holes (Colbert and Mushotzky, 1999; Miller, 2006; Maccarone et al., 2007; Koliopanos et al., 2017). However, the discovery of coherent X-ray pulsations in M82 X-2 (Bachetti and others, 2014), which exhibits an extraordinary luminosity of $L_{\mathrm{X}}\simeq 1.4\times 10^{40}\,\rm erg\,s^{-1}$} and a pulse period of $1.37\,\rm s$, provided compelling evidence that some ULXs are powered by accreting pulsars rather than black holes. In these systems, matter channeled by the strong magnetic field onto the stellar surface forms hot spots that produce pulsed X-ray emission as the star rotates.
@@
Since the discovery of M82 X-2, numerous other ULX pulsars (ULXPs) have been identified, and their observed properties are summarized in Table 1. Of particular interest is the possible detection of a narrow cyclotron scattering feature at $\sim 4.5\,\rm keV$ in M51 ULX-8 (Brightman et al., 2018). Although pulsations have not been detected from this source, the feature has been interpreted as a proton cyclotron line, implying a surface multipole magnetic field as strong as $B\sim 10^{15}\,\rm G$. The X-ray luminosities of ULXPs are typically in the range of $10^{39}-10^{41}\,\rm erg\,s^{-1}$, which is above the Eddington luminosity for accreting pulsars, $L_{\rm Edd}\approx 1.8\times 10^{38}\,(M/M_{\odot})\,\rm erg\,s^{-1}$, where $M$ is the mass of the compact object. This challenges our understanding of accretion physics under extreme magnetic fields, high temperatures, and super-Eddington regimes, making ULXPs unique laboratories for exploring high-energy astrophysical processes.
@@
The origin of the extremely high luminosity observed in ULXPs remains a subject of ongoing debate. Proposed models for ULXPs span a broad range of surface magnetic field strengths. At relatively low fields ($\sim 10^{11}\,\rm G$), the observed luminosities may be explained by strong radiatively driven outflows and geometrical beaming (Middleton and King, 2016, 2017; King and Lasota, 2019; Lasota and King, 2023). At the other extreme, magnetar-like fields ($\sim 10^{14}\,\rm G$) can suppress radiation pressure (Basko and Sunyaev, 1976; Tong, 2015; Dall’Osso et al., 2016; Ekşi et al., 2015; Chandra, 2026) and confine accretion columns near the magnetic poles (Mushtukov et al., 2015; Inoue, 2020). Beyond the field strength itself, several additional factors may substantially influence the accretion dynamics in ULXPs, including complex, non-dipolar magnetic field configurations (Israel and others, 2017a; Tsygankov et al., 2018), enhanced neutrino cooling (Mushtukov et al., 2018; Asthana et al., 2023), and photon-bubble instabilities (Arons, 1992; Begelman, 2006; Zhang et al., 2021).
@@
On the other hand, three primary mechanisms can contribute to neutrino emission in X-ray binaries: (i) high-energy protons accelerated within the system, such as in jets, relativistic winds, or pulsar magnetospheres. These protons undergo collisions that predominantly produce pions, which then decay into gamma rays and neutrinos (Neronov and Ribordy, 2009; Sahakyan et al., 2014; Kosmas et al., 2023; Christiansen et al., 2006; Ducci et al., 2025); (ii) proton-photon interactions in the magnetosphere of the accreting pulsar (Bednarek, 2009); (iii) electron-positron pairs of the accretion column of the bright X-ray pulsars, where pair annihilation leads to neutrino emission (Mushtukov et al., 2018, 2019; Asthana et al., 2023; Mushtukov et al., 2025).
@@
In previous studies of electron-positron pair annihilation in the accretion columns of bright X-ray pulsars and ULXPs, however, the properties of the thermal mound, as well as the structure and equation of state (EOS) of the dense matter inside pulsars, were often neglected (Mushtukov et al., 2018; Asthana et al., 2023). In fact, the EOS of supranuclear matter inside compact objects remains one of the most challenging problems in both microphysics and astrophysics, hopefully to be solved with the multi-messenger astronomy (Lattimer and Prakash, 2004; Baiotti, 2019). Due to the non-perturbative nature of the fundamental strong interaction at low energies, the EOS of dense matter at nuclear densities is still poorly understood, nevertheless, strangeness should play an important role in this energy scale (Lai et al., 2023). Witten (1984) proposed that the true ground state of dense matter might be quark matter, composed of nearly free $u$, $d$, and $s$ quarks. This concept suggests that compact objects resembling pulsars could be quark stars (QSs) rather than traditional neutron stars (NSs), as outlined in the MIT bag model in the asymptotic freedom regime (Alcock et al., 1986). Xu (2003) proposed, however, that building units of supranuclear matter could instead be strange quark clusters (renamed “strangeons”), rather than itinerating quarks, due to the fact that the coupling between them is still very strong (Lai and Xu, 2017). These strangeons, formed by bound $u$, $d$, and $s$ quarks, represent a unique state of matter in which quarks condense in position space rather than momentum space. The term strangeon stars (SSs) was coined to describe such compact objects (Xu and Guo, 2017; Lai and Xu, 2017; Li et al., 2022; Yuan et al., 2025).
@@
Strangeon matter, like strange quark matter, consists of nearly equal numbers of $u$, $d$, and $s$ quarks. However, in contrast to strange quark matter, quarks in strangeon matter are localized inside strangeons due to the strong quark-quark coupling. NSs, QSs, and SSs share similarities but also differ significantly. In SSs, quarks are localized in strangeons, much like neutrons in NSs. However, unlike neutrons, a strangeon can contain more than three valence quarks, restoring light-flavour symmetry. Additionally, SSs are self-bound by the strong force, with their surface matter also composed of strangeons, similar to QSs (Xu, 2003).
@@
SSs can explain many key observational phenomena in astrophysics. The EOS of SSs is sufficiently stiff to account for the observed masses of pulsars (Demorest et al., 2010; Antoniadis and others, 2013), while pulsar glitches may be attributed to starquakes (Zhou et al., 2014). Furthermore, SSs are capable of explaining X-ray flares and bursts in magnetar candidates (Xu et al., 2006), the plateau phase in gamma-ray bursts (Dai et al., 2011), the quasi-periodic oscillations in SGR 1806-20 (Li et al., 2023), and shaping the energy budget of various power sources in the context of magnetar starquake triggering mechanisms (Wang et al., 2024). Additionally, the strangeness barrier plays a crucial role in understanding Type I X-ray bursters (Li et al., 2015) and X-ray-dim isolated neutron stars (Wang et al., 2017, 2018).
@@
In this work, we investigate the thermal mound properties within the SS framework (Xu, 2003) and estimate the neutrino luminosity relevant to ULXPs. A notable feature of the polar mound of SSs is the sharp discontinuity in mass density, resulting from both the Coulomb (Xu and Qiao, 1999; Hu and Xu, 2002; Xu, 2014) and the strangeness (Xu, 2014; Li et al., 2015; Wang et al., 2017, 2018) barriers. Most of the total gravitational energy of accretion is stored in the matter above the surface of density discontinuity on the polar cap of an SS, whereas an NS does not exhibit such a discontinuity. This therefore results in a lower effective heat capacity for SS and a higher one for NS, and consequently a high temperature and large neutrino emissivity of SS. The detailed quantitative calculations in the following sections demonstrate consistency with these qualitative results.
@@
NGC 4559 X7 2.6 1-1 55 7.5 [14]
  • [1][{1}] Bachetti and others (2014);
  • [2][{2}] Israel and others (2017a); Fuerst and others (2023);
  • [3][{3}] Carpano et al. (2018); Vasilopoulos et al. (2018); Walton and others (2018);
  • [4][{4}] Sathyaprakash and others (2019);
  • [5][{5}] Trudolyubov et al. (2007);
  • [6][{6}] Motch et al. (2013); Fürst and others (2016); Israel and others (2017b); Fuerst and others (2018);
  • [7][{7}] Quintin et al. (2021);
  • [8][{8}] Pintore and others (2021, 2025);
  • [9][{9}] Wilson-Hodge and others (2018);
  • [10][{10}] Weng et al. (2017);
  • [11][{11}] Vasilopoulos and others (2020);
  • [12][{12}] Rodríguez Castillo and others (2020); Vasilopoulos et al. (2020);
  • [13][{13}] Brightman et al. (2018); Bachetti and others (2022);
  • [14][{14}] Pintore et al. (2025).
@@ 2.1 Dynamical structure of the accretion column
Figure 1: Schematic depiction of the magnetic polar regions with radius $R_{\rm s}\simeq 10^{5}\,\rm cm$ and height $H_{\rm mound}$ of polar mound in an X-ray pulsar. $\Theta$ is the opening angle of the magnetic funnel. This diagram is adapted from Inoue (2020).
@@
Radiation pressure plays a crucial role in determining the dynamical structure of accretion flows in high-luminosity X-ray pulsars. At sufficiently high accretion rates, a shock forms above the stellar surface (Inoue, 1975; Basko and Sunyaev, 1976; Inoue, 2020). This shock must be radiative in nature to effectively dissipate the kinetic energy of the infalling matter (Becker, 1998). Models of such accretion columns are essential for interpreting observed radiation luminosities (Mushtukov et al., 2015), pulse profiles (Klochkov et al., 2008; Becker et al., 2012), continuum spectra (Becker and Wolff, 2006), cyclotron line spectra (Staubert and others, 2019), and polarization observations of high-luminosity X-ray pulsars in high-mass X-ray binaries. Moreover, the accretion column exhibit quasi-periodic oscillations using the magnetohydrodynamics simulations (Zhang et al., 2022, 2023, 2025)
@@
In highly magnetized accreting pulsars, the plasma is funneled by the strong dipolar magnetic field toward the magnetic polar caps, forming narrow accretion columns above the stellar surface. In these regions, the inflowing matter is nearly constrained to move along magnetic field lines, and the lateral motion can be neglected compared to the radial inflow. Assuming one-dimensional, steady-state accretion, the structure of the accretion column is determined by the conservation of mass, momentum, and energy (Inoue, 1975; Basko and Sunyaev, 1976; Becker, 1998).
Following the variational derivation of Inoue (1975), the dynamical equations describing the vertical structure of the magnetic polar regions can be written as (see also Inoue, 2020),
@@
$r_{\rm D}\equiv\frac{3\kappa_{\rm Th}\dot{M}}{8\pi c}=1.4\times 10^{5}\left(\frac{\dot{M}}{10^{17}\rm{g\;s}^{-1}}\right)\;\rm{cm}\,$, (3)
where $\kappa_{\rm Th}$ is the opacity of the Thomoson scattering in the tangential direction of the polar cone, $\dot{M}$ is the mass accretion rate onto the stellar surface. The quantity $r_{\rm D}$ represents the distance that the infalling matter advances within the time it takes for a photon to diffuse from the central axis to the surface of the cone. Eq. (1) describes the evolution of the energy density, balancing gravitational acceleration $(GM/r^{2})$ and radiative dissipation. Eq. (2) governs the density stratification, where the compressive effects of gravity compete with radiative pressure support. Together, these equations determine the vertical structure of the accretion column and provide the foundation for understanding the formation of the thermal mound and the associated energy transport in the polar regions.
We adopt the same boundary conditions as in Inoue (2020). At the top boundary of the polar cone, we set the mass density to
@@
In Fig. 2, we show the radiation energy, mass density, pressure, velocity, optical depth, and temperature as functions of the column height for different mass accretion rates. The optical depth $\tau$ of the Thomson scattering in the tangential direction of the polar cone is given by (Inoue, 2020)
$\tau=\kappa_{\rm Th}\rho r\Theta_{0}\,.$ (7)
Here, the opening angle $\Theta_{0}$ of the magnetic polar region can be written as (Inoue, 2020)
$\Theta_{0}=$  $\,4\times 10^{-2}\,\left(\frac{R}{10^{6}\,\rm cm}\right)^{1/2}\left(\frac{\dot{M}}{10^{17}\,\rm g\,s^{-1}}\right)^{1/7}\left(\frac{M}{M_{\odot}}\right)^{1/14}$
$\,\left(\frac{\mu_{\rm M}}{10^{30}\,\rm G\,cm^{3}}\right)^{-2/7}\,$, (8)
where $\mu_{\rm M}$ is the magnetic dipole moment of the pulsar. It can be observed that an extremely high mass accretion rate results in a taller accretion column. Moreover, the pressure and temperature at the base of the column remain independent of the accretion rate because the magnetic pressure counterbalances the radiation pressure.
In particular, the temperature in the polar mound exceeds $10^{9}\,\rm K$, leading to energy losses through neutrino emission (Beaudet et al., 1967; Kaminker et al., 1994). The primary mechanism for neutrino production is the annihilation of electron-positron pairs, $e^{-}+e^{+}\rightarrow\nu+\bar{\nu}$ (Beaudet et al., 1967). At high temperature regime the pair annihilation remains the dominant process for neutrino production with the corresponding energy losses of $Q_{\nu}\approx 4\times 10^{24}\,\left(\frac{T}{10^{10}\,\rm K}\right)^{9}\,{\rm erg\,cm^{-3}\,s^{-1}}$.
Figure 2: Profiles of physical quantities within the magnetic polar regions as functions of height, with different colors indicating different mass accretion rates. We adopt a stellar model with $M=1.4\,M_{\odot}$ and $R=10^{6}\,\rm cm$. The first column presents the radiation energy density, mass density, and pressure, while the second column shows the corresponding velocity, optical depth, and temperature profiles. Normalization parameters are defined as $\varepsilon_{0}=GM/R$, $\rho_{0}=3B^{2}R/(8\pi GM)\sim 10^{2}\,\rm g\,cm^{-3}$, and $P_{0}=B^{2}/(8\pi)$, assuming a surface magnetic field strength of $B=10^{12}\,\rm G$. The characteristic temperature, $T_{\rm C}$, is calculated via $T_{\rm C}=(\rho\varepsilon/a)^{1/4}$, where $a$ is the radiation constant. The height of the polar mound $H_{\rm mound}\approx 0.8\,\rm km$.